OpenStudy (anonymous):

prove/disprove a prime number p can't be written as sum of 3 squares .

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ 6^2+6^2+1^2=73 }\end{align}\)

OpenStudy (anonymous):

hmm counter example :|

OpenStudy (anonymous):

that works actully

OpenStudy (mathmath333):

simplest example \(\large\tt \begin{align} \color{black}{ 1^2+1^2+1^2=3 }\end{align}\)

ganeshie8 (ganeshie8):

you want to find a prime that cannot be written as sum of squares of 3 integers ?

OpenStudy (anonymous):

i just wanted to be confident as much as i can about primes properties hmm at which type of primes we can write them as sum of 3 squares :| idk if its even possible i wanna any integers

ganeshie8 (ganeshie8):

you cannot write primes of form 8k-1 as sum of squares of 3 integers for example : 23 cannot be written as sum of squares of 3 integers

OpenStudy (anonymous):

what about others?

OpenStudy (anonymous):

like what about 8k+7

ganeshie8 (ganeshie8):

thats same as 8k-1

OpenStudy (anonymous):

lol ithought that was 8k+1 :(

ganeshie8 (ganeshie8):

in mod8 primes can only be of form 8k+1 8k+3 8k+5 8k+7

ganeshie8 (ganeshie8):

you know that primes of form 8k+7 cannot be expressed as sum of 3 squares you want to analyze the remaining ?

OpenStudy (anonymous):

ok i got now why i cant

OpenStudy (anonymous):

yes sure

ganeshie8 (ganeshie8):

why ?

OpenStudy (anonymous):

8k+1=2(2k+1)^2-(8k^2+1) ?

ganeshie8 (ganeshie8):

so ?

OpenStudy (anonymous):

xD ok show me what got

ganeshie8 (ganeshie8):

Notice a square only leaves remainders : {0, 1, 4} in mod8

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

taking any combination of them can never add up to 7 \(\blacksquare \)

OpenStudy (anonymous):

i see

ganeshie8 (ganeshie8):

0 + 0 + 0 = 0 mod 8 1 + 1 + 1 = 3 mod 8 4 + 4 + 4 = 4 mod 8 0 + 1 + 1 = 2 mod 8 0 + 1 + 4 = 5 mod 8 ...

ganeshie8 (ganeshie8):

it is never 7 mod 8

ganeshie8 (ganeshie8):

infact that proves that no integer of form 8k+7 can be expressed as sum of 3 squares see this http://math.stackexchange.com/questions/779784/show-that-an-integer-of-the-form-8k-7-cannot-be-written-as-the-sum-of-three

OpenStudy (anonymous):

ok got it